OpenStudy (anonymous):

Is it possible to simplify (-32)^(3/5), if so, what does it simplify to?

OpenStudy (zzr0ck3r):

\((-32)^{\frac{3}{5}}=((-32)^{\frac{1}{5}})^3=(-2)^3=-8\)

OpenStudy (anonymous):

thanks, so you could do it even if it's a negative number?

OpenStudy (anonymous):

That would depend on the number in the denominator. If it's even, you can't get a real root.

OpenStudy (zzr0ck3r):

are you in R?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

but anyways, thanks guys!

OpenStudy (zzr0ck3r):

are you dealing with complex numbers or just real numbers?

OpenStudy (anonymous):

idk...

OpenStudy (anonymous):

real.

OpenStudy (zzr0ck3r):

ok you would, so its real...

OpenStudy (zzr0ck3r):

ok all good

OpenStudy (anonymous):

thanks!